OpenStudy (anonymous):

steps please... (11x+12)=3(x+12) answer is 3

OpenStudy (anonymous):

(11x+12)=3(x+12) now you multiply 3 times x and 12 (11x+12)=(3x+36) subtract 3x from 11x and 12 from 36 8x=24 now isolate the x by dividing the 24 by 8 x=3

OpenStudy (anonymous):

Thanks alot

OpenStudy (anonymous):

with pleasure